Multiple Versus Single Overhead Rates, Activity Drivers

Deoro Company has identified the following overhead activities, costs, and activity drivers for the coming year:

Activity Expected Cost Activity Driver Activity Capacity
Setting up equipment $497,860 Number of setups 620
Ordering costs 366,000 Number of orders 18,300
Machine costs 856,000 Machine hours 42,800
Receiving 432,600 Receiving hours 10,300

Deoro produces two models of dishwashers with the following expected prime costs and activity demands:

Model A Model B
Direct materials $605,000 $773,000
Direct labor $482,000 $477,000
Units completed 15,300 7,600
Direct labor hours 6,000 1,700
Number of setups 420 200
Number of orders 5,700 12,600
Machine hours 24,700 18,100
Receiving hours 2,600 7,700

The company's normal activity is 7,700 direct labor hours.

Required:

1. Determine the unit cost for each model using direct labor hours to apply overhead. Round intermediate calculations and final answers to nearest cent.

Unit Cost
Model A $
Model B $

2. Determine the unit cost for each model using the four activity drivers. Round your answers to nearest cent.

Unit Cost
Model A $
Model B $
Answer #1
Overhead rate = Total overheads/Total labor hours              279.54 per hour
Model A Model B
Direct Material            605,000            773,000
Direct Labor            482,000            477,000
Overheads        1,677,240            475,218
Total cost        2,764,240        1,725,218
Number of units              15,300                7,600
Unit cost              180.67              227.00
2.
Activity Expected Cost Activity Driver Activity Capacity Activity Rate
Setting up equipment 497860 Number of setups 620 803
Ordering costs 366,000 Number of orders 18,300 20
Machine costs 856,000 Machine hours 42,800 20
Receiving 432,600 Receiving hours 10,300 42
Total overheads        2,152,460
Model A Model B
Direct Material            605,000            773,000
Direct Labor            482,000            477,000
Overheads
Setting up equipment            337,260            160,600
Ordering costs            114,000            252,000
Machine costs            494,000            362,000
Receiving            109,200            323,400
Total cost        2,141,460        2,348,000
Number of units              15,300                7,600
Unit cost              139.96              308.95

Model A 605000 482000 =279.54*6000 =SUM(B3:B5) 15300 =B6/B7 Model B 773000 477000 =279.54*1700 =SUM(C3:C5) 7600 =C6/07 2 3 Di